Celia Mae starts a ball rolling up a 4.5° hill on a planet where g is half of that on Earth with a velocity of 3.2 m/s. Once she starts it moving, she just watches it roll.

A. How far does the ball roll up the incline before it stops?
B. How long does it take the ball to stop?
C. How far does the ball roll up the incline in 2.5 seconds?
1
Expert's answer
2019-09-23T09:10:58-0400

A. From the conservation of energy:


"0.5mv^2=m(0.5g)s\sin{4.5\degree}"

"3.2^2=9.8\sin{4.5\degree}s"


"s=13.3\ m"

B.


"t=\frac{v}{0.5g\sin{4.5\degree}}=\frac{2(3.2)}{9.8\sin{4.5\degree}}=8.3\ s"

C.


"d=vT-0.5(0.5g)\sin{4.5\degree}T^2"

"d=(3.2)(2.5)-0.25(9.8)\sin{4.5\degree}2.5^2=5.6\ m"